[figure: 9.7 cm]

DETERMINING THE FERMI SURFACE FROM ANGLE RESOLVED PHOTOEMISSION SPECTROSCOPY (ARPES)

Konrad Matho Centre de Recherches sur les Très Basses Températures, laboratoire associé à l’Université Joseph Fourier, CNRS, BP 166, 38042 Grenoble Cedex 9, France

A point k_{F} on the Fermi surface is the locus of a steplike feature in the momentum distribution n_{k}. Energy integrated ARPES spectra give experimental access to n_{k}(T) for strongly correlated electrons. Various modified criteria for k_{F} have been proposed, taking into account that n_{k} is discontinuous only in the zero temperature limit. In the absence of particle hole symmetry around the chemical potential, I find, that k_{F} is best characterized by the stability of n_{k}(T) ≈ n_{k}(O). The actual value may be ≠ 1/2. The “inflexion-point”-criterion is less stable. Acknowledging the help of several poles (in the complex plane), I am happy to present my conclusions at a conference in their homeland.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : oral

Corresponding author : Konrad Matho

Address for correspondence : CRTBT-CNRS BP 166 38042 Grenoble Cedex France

Email address : matho@labs.polycnrs-gre.fr